A flywheel of mass 4kg and radius 0.5m rotates freely on a horizontal axis.a block of mass 2.0 kg hangs by a string that is tightly wrapped around the flywheel. Find the angular velocity and the speed of the block when the block has fallen through 1.6m
1
Expert's answer
2021-07-19T09:51:07-0400
"mgh=0.5mv^2+0.5I\omega^2\\mgh=0.5mv^2+0.5(0.5Mv^2)
\\(2)(9.8)(1.6)=0.5(2)v^2+0.25(4)(v^2)\\v=4.0\frac{m}{s}\\\omega=\frac{4.0}{0.5}=8.0\frac{rad}{s}"